1. Given:AB=(3x-5)cm,BC=(2y-7)cm,CD=(X+7)cm and AD=(y+3)cm.
a.what is the value of x?
b.how long is AB?
c.what is the value of y?
d.how long is AD?
e.how did you solve for the values of x and y?
f.what property did you apply to determine the lengths of AB and AD?
2.BAD measures (2a+25)° while BCD measures (3a-15)°
a.what is the value of a?
b.what is the m BAD
c.what is the m CBA?
e.how did you find the value of a?
f.what property did you apply to solve for m CBA?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

1. Given:AB=(3x-5)cm,BC=(2y-7)cm,CD=(X+7)cm and AD=(y+3)cm.
AB=CD, BC =AD opposite sides of parallelogram
AB=(3x-5)cm,BC=(2y-7)cm,CD=(X+7)cm and AD=(y+3)cm.
(3x-5)=(X+7) .
2x=12
x=6
a.what is the value of x? x=6
b.how long is AB?
AB=3x-5 = 13
c.what is the value of y?
2y-7 =y+3
y=10
d.how long is AD? y+3 . AD=13
e.how did you solve for the values of x and y?
f.what property did you apply to determine the lengths of AB and AD?
AB=CD, BC =AD opposite sides of parallelogram
2.BAD measures (2a+25)° while BCD measures (3a-15)°
(2a+25)° =(3a-15)°
a=40 deg
a.what is the value of a? 40 deg
b.what is the m BAD = 2a +25 =105 deg
c.what is the m CBA?
m CBA = 75 deg
e.how did you find the value of a?
mBAD + mA CBA=180 deg

f.what property did you apply to solve for m CBA?
Adjacent angles of a parallelogram are supplementary

Answer by ikleyn(53765)

.

As your post is worded, printed, posted and presented, it MAKES no SENSE, since the shape ABCD is not defined in it.

In normal problem,  it  MUST  BE  DEFINED  in the very first sentence.

If it is not defined in the very first sentence,  it means that the entire post is  DEFECTIVE.